Spectral widths of H+2 multiphoton dissociation with short intense laser pulses.

Atabek O. and Jolicard G.

Abstract

Photodissociation spectra of H_{2}, obtained using ultrashort (a few femtoseconds), intense (whitin the range of 1012 to 1014 W/cm^{2}) laser pulses (with a peak frequency of 3.761 eV), and displaying wellseparate proton kinetic-energy peaks corresponding to one-, two-, or three-photon absorption are analyzed. Whitin some approximations, the validity range of which is related to the aforementioned assumptions on the laser and spectrum characteristics, a simple analytic formula is deduced. It gives the spectral widths of the observed peaks in terms of the total number of photons which are absorbed and/or emitted during the dissociation process. A quantitative support to some recent dynamical intepretations is provided. The successive absorption of three photons seems basically to be the initial step which, at intermediate intensities, leads to the three-photon peaks. A subsequent emission of one (or two) photons is responsible for the two-(or one-) photon peaks. Bond softening occurs at higher intensities by the lowering of potential- energy barriers. Dissociation proceeds through these barriers by reflection and tunneling. It involves at least three photons to end in a one-photon peak.